OpenStudy (anonymous):

Please help! Prove - tan2x + sec2x = 1

OpenStudy (anonymous):

Rewrite everything in terms of sine and cosine: \[-\frac{\sin^2 x}{\cos^2 x}+\frac{1}{\cos^2 x} = 1\] Then just add the fractions, and use the Pythagorean identity to show that it's true: \[\frac{1-\sin^2 x}{\cos^2 x} = \frac{cos^2 x}{\cos^2 x} = 1\]

OpenStudy (jdoe0001):

http://www.gradeamathhelp.com/image-files/trigonometric-identities-pythagorean.gif <--- notice the 2nd identity, solve for "1"